# TopER: Topological Embeddings in Graph Representation Learning

### v2

@@ Appendix C Proofs of Stability Theorems
Lemma 1.
Skraba and Turner, (2020) Let $\mathcal{X}$ be a compact metric space, and $f,g:\mathcal{X}\to\mathbb{R}$ be two filtration functions. Then, for any $p\geq 1$, we have $\mathcal{W}_{p}(\rm{PD}_{k}(\mathcal{X},f),\rm{PD}_{k}(\mathcal{X},g))\leq\|f-g\|_{p}$
The next lemma is on the stability of Betti curves by Dłotko and Gurnari, (2023) [Proposition 1].
Lemma 2.
Dłotko and Gurnari, (2023) Let $\beta_{k}(\mathcal{X})$ is the $k^{th}$ Betti function obtained from the persistence module $\rm{PM}_{k}(\mathcal{X})$.
@@
Theorem 2.
@@
First, we employ the stability of Betti curves by Lemma 2.
@@
Hence to obtain $\rm{TE}_{f}(\mathcal{X})=(a_{f},b_{f})$, we fit least squares line $y=a_{f}+b_{f}x$ to the set of $N$ points in $\mathbb{R}^{2}$, i.e., $\mathcal{Z}_{f}=\{(\beta^{f}_{0}(\epsilon_{i}),\beta^{f}_{1}(\epsilon_{i}))\}_{i=1}^{N}$. Similarly, we obtain $\rm{TE}_{g}(\mathcal{X})=(a_{g},b_{g})$ by fitting least squares line to $\mathcal{Z}_{g}=\{(\beta^{g}_{0}(\epsilon_{i}),\beta^{g}_{1}(\epsilon_{i}))\}_{i=1}^{N}$. By Equation 1, we have
@@
Now, by the stability of least squares fit with respect to Hausdorff distance (Chernov et al., (2012) [Theorem 3.1]), we have
@@
Hence, when we combine Equations 2 and 3, we have
@@
By combining the above result with Lemma 1, we obtain the following corollary.
Corollary 2.
@@
By combining Equations 4 and 5, we conclude

### v3

@@ Appendix C Proofs of Stability Theorems
Lemma˜3.6. [66] Let $\mathcal{X}$ be a compact metric space, and $f,g:\mathcal{X}\to\mathbb{R}$ be two filtration functions. Then, for any $p\geq 1$, we have $\mathcal{W}_{p}(\rm{PD}_{k}(\mathcal{X},f),\rm{PD}_{k}(\mathcal{X},g))\leq\|f-g\|_{p}$
The next lemma is on the stability of Betti curves by [26] [Proposition 1].
Lemma˜3.7. [26] Let $\beta_{k}(\mathcal{X})$ is the $k^{th}$ Betti function obtained from the persistence module $\rm{PM}_{k}(\mathcal{X})$.
@@
Theorem˜3.4 Let $\mathcal{X}$ be a compact metric space, and $f,g:\mathcal{X}\to\mathbb{R}$ be two filtration functions. Then, for some $C>0$,
@@
First, we employ the stability of Betti curves by Lemma˜3.7.
@@
Hence to obtain $\rm{TE}_{f}(\mathcal{X})=(a_{f},b_{f})$, we fit least squares line $y=a_{f}+b_{f}x$ to the set of $N$ points in $\mathbb{R}^{2}$, i.e., $\mathcal{Z}_{f}=\{(\beta^{f}_{0}(\epsilon_{i}),\beta^{f}_{1}(\epsilon_{i}))\}_{i=1}^{N}$. Similarly, we obtain $\rm{TE}_{g}(\mathcal{X})=(a_{g},b_{g})$ by fitting least squares line to $\mathcal{Z}_{g}=\{(\beta^{g}_{0}(\epsilon_{i}),\beta^{g}_{1}(\epsilon_{i}))\}_{i=1}^{N}$. By Equation˜1, we have
@@
Now, by the stability of least squares fit with respect to Hausdorff distance ([16] [Theorem 3.1]), we have
@@
Hence, when we combine Equations˜2 and 3, we have
@@
By combining the above result with Lemma˜3.6, we obtain the following corollary.
Corollary˜3.5 Let $\mathcal{X}$ be a compact metric space, and $f,g:\mathcal{X}\to\mathbb{R}$ be two filtration functions. Then, for some $C>0$,
@@
By combining Equations˜4 and 5, we conclude
@@
Appendix D Synthetic Experiments
This section describes experiments performed on the Erdos-Renyi synthetic graph. We conducted two experiments by applying TopER on this graph and Principal Component Analysis (PCA). The goal is to compare the performance, runtime, and interpretability of the two models. From the plots shown in Figures 8, 9, and 10, we see that TopER is interpretable compared to when PCA is used.
Figure 8: TopER plots showing pivot vs growth for each function—degree centrality, closeness, degree, and popularity when threshold=50.
Figure 9: Persistent Homology PCA plots embeddings (Betti 0 and Betti 1) showing component 1 vs component 2 for each function—closeness and degree when threshold=50 and the respective filtrations —sublevel and superlevel.
Figure 10: Persistent Homology PCA plots embeddings (Betti 0 and Betti 1) showing component 1 vs component 2 for each function—degree centrality and popularity when threshold=50 and the respective filtrations —sublevel and superlevel.
Appendix E Broader Impact
This work advances the field of graph representation learning by introducing a topological approach that is both interpretable and scalable. By leveraging the structural insights of persistent homology without incurring its prohibitive computational costs, TopER enables more efficient and insightful analysis of complex graph-structured data. This has the potential to benefit a range of scientific and industrial domains where graph data is prevalent, including bioinformatics, social network analysis, and infrastructure monitoring. In particular, the ability to generate low-dimensional and interpretable embeddings could assist researchers in visual analytics, pattern discovery, and model debugging. At the same time, we acknowledge that the use of graph representations—especially in social networks or biological datasets—may carry ethical concerns around data privacy, representational bias, or unintended consequences of automated decision-making. While TopER itself is an unsupervised and domain-agnostic method, its application must be governed by domain-specific ethical considerations. To support responsible use, we emphasize interpretability and transparency in our design, and we release our code and visualizations to promote reproducibility and community oversight.